OpenStudy (anonymous):

The answer is approximately 22, but I can't figure out how to get there. 33=31.5+1.1log(x+1)

OpenStudy (solomonzelman):

\[33=31.5+1.1\log(x+1)\]subtract 31.5 from both sides, \[1.5=1.1~\log(x+1)\]divide both sides by 1.1\[1.36=\log(x+1)\] can you take it from here ?

OpenStudy (anonymous):

That's actually the part I'm stuck on. Am I just suppose to take the log of both sides?

OpenStudy (solomonzelman):

\[1.36=\log(x+1)\]unspecified base,\[1.36=\log_{10}(x+1)\] Now, focus and follow.... \[1.36=\log_{10}(x+1)\]\[1.36 \times \log_{10}10=\log_{10}(x+1)\]\[ \log_{10}10^{1.36}=\log_{10}(x+1)\]\[10^{1.36}=(x+1)\] can you do it from here ?

OpenStudy (anonymous):

Ah, I see now. I can take it from there. Thank you so much!

OpenStudy (solomonzelman):

You are welcome, always remember this tip, of using \[\log_xx~~(The~~magic~~1)\]